# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 1. Introduction
The focus of this paper is the regularity of two dimensional Lipschitz solutions to the elliptic equation:
$\operatorname{div}(G(Du))=0\text{ in }\mathcal{D}^{\prime}(B_{1}).$ (1.1)
To motivate the following discussion, let us consider, for the moment, its variational counterpart:
$\operatorname{div}(Df(Du))=0\text{ in }\mathcal{D}^{\prime}(B_{1}),$ (1.2)
which is the Euler-Lagrange equation satisfied by minima of the functional
@@
It is classical that solutions $u$ to (1.2) are smooth provided $f$ is convex, $C^{2}(\mathbb{R}^{2})$, and
@@
In higher dimensions, this is due to the celebrated theorems of De Giorgi-Nash-Moser [16, 48, 45], while in two dimensions it has been known since the work of Morrey [44]. In fact, in two dimensions, we will see below that one may assume only one of the bounds in (1.4) and still infer $C^{1}$ regularity of Lipschitz solutions to (1.2), provided $f$ is strictly convex. In the context of the more general equation (1.1), strict convexity of $f\in C^{1}(\mathbb{R}^{2})$ should be replaced by strict monotonicity of $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$:
$(G(b)-G(a),b-a)\geq\sigma(|b-a|),\quad\forall a,b\in\mathbb{R}^{2},$ (1.5)
@@
$\sigma:[0,+\infty)\to[0,+\infty)\text{ is a strictly increasing, convex function.}$ (1.6)
In the following, for brevity we will often say that $G$ fulfills (1.5) to say that $G$ fulfills (1.5) for $\sigma$ fulfilling (1.6). This discussion motivates the following question, which is the starting point of this paper. Assume $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ and strictly monotone:
are Lipschitz solutions to (1.1) $C^{1}$, or at least partially regular? (Q)
One motivation for this question is that many important and natural energies, such as the $p$-Laplacian, do not enjoy the strong ellipticity property (1.4) and their regularity theory is more complicated, even in two dimensions. See [42] for an extensive review of problems where (1.4) fails for $|\xi|\to 0$ and $|\xi|\to\infty$. Cases where (1.4) fails in a more general fashion are much less understood, but have seen a lot of progress in recent years, starting with the work of D. De Silva and O. Savin [20]. In order to review the literature, we introduce some (loosely defined) notation. We let $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ be the sets in $B_{\operatorname{Lip}(u)}$ where the symmetric part of $(DG)^{-1}$, respectively of $DG$, has a zero eigenvalue, see (1.16) for the formal definition. In the variational case (1.2), $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ should be thought as the sets of points in $B_{\operatorname{Lip}(u)}$ at which $D^{2}f$ has an infinite eigenvalue and a zero eigenvalue, respectively. Then, the answer to (Q) is yes, provided:
  • $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ is empty or $\mathcal{D}_{-}$ is finite, [20];
  • $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ is finite, [35];
  • $G$ is $\delta$-monotone, see [2, Chapter 16.4.1]. In this case, $u$ needs not be Lipschitz;
  • DuDu lies in a curve and it solves (1.1)-(1.5) for $\sigma(t)\sim t^{4}$, [37].
Partial regularity is also obtained in [20] for an obstacle problem involving minimizers of (1.3) under the constraint that $Du$ lies in a convex polygon, with a degeneracy set $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ that may contain the polygon’s boundary. Further, an example given in [35, Theorem 1.5] shows that, under the mere assumption (1.5), Lipschitz solutions of (1.1) may have point singularities, hence one cannot expect better in general than partial regularity in (Q).
Having asked that $u\in\operatorname{Lip}(B_{1})$, (Q) becomes a question about controlling the oscillations in the gradient of $Du$. Of course, a similar question could be asked about concentration effects, namely starting from a $W^{1,p}(B_{1})$ function. Some results in this direction, and valid in all dimensions, can be found in [14, 15] and references therein. Notice, however, that in the unbounded gradient case, further and careful assumptions need to be placed on $u$, $f$ or $G$, as counterexamples to regularity of very weak solutions to elliptic equations show, see [9, 30]. Besides, in two dimensions the Lipschitz property of $u$ can be inferred naturally in many situations, compare [24, §12.4] and [8], valid in all dimensions. Another direction in which to extend this analysis is, of course, higher dimensions, see [43, 41, 40]. In that case, however, the picture is much less clear than in two dimensions: to the best of our knowledge, in that context even partial regularity for $u$ solving (1.2) for $f$ strictly convex, smooth, and fulfilling only the upper bound in (1.4) is unknown.
As said, the aim of this work is to answer Question (Q). Let us give a rough idea of our main results, deferring a more precise description to §1.2-1.3-1.4 below. The starting point is to write, formally, (1.1) as
$\langle DG^{s}(Du),D^{2}u\rangle=0.$
Then, if $Du(x)\notin\mathcal{D}_{+}\cap\mathcal{D}_{-}$, we can expect some sort of ellipticity in a neighborhood of $x$, while all ellipticity is lost if $Du(x)\in\mathcal{D}_{+}\cap\mathcal{D}_{-}$. However, if $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ has a sufficiently nice structure, we can interpret
$Du\in\mathcal{D}_{+}\cap\mathcal{D}_{-}$ (1.7)
as an additional equation, coupled with (1.1). Our results show (partial) $C^{1}$ regularity under, for instance, the assumption that $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ is a finite union of curves. In that case, (1.7) coupled with (1.1) becomes a hyperbolic problem: to study it, we rely on the theories of viscosity solutions to Hamilton-Jacobi equations and entropy solutions of scalar conservation laws. It is important to note that our results recover and extend all the results contained in the previous list where, essentially, $\mathcal{D}_{+}\cap\mathcal{D}_{-}$ had to be taken of dimension zero.
To implement the strategy outlined above, the crucial point is to obtain a good localization property of the gradient of a solution of (1.1), Theorem A. This property, inspired by the beautiful separation argument of [22], tells us that oscillations of the gradient are small in $L^{\infty}$ in any small enough neighborhood of a differentiability point (where oscillations would a priori be small only in $L^{p}$ for $p<\infty$). Through Theorem A, we are able to efficiently split $B_{1}$ into regular and singular points, based on the Lipschitz blow-ups of $u$. Our next result, Theorem B, shows that the singular set is $\mathcal{H}^{1}$-negligible. Next, in Theorem C, we show that one can associate to a solution to (1.1) a measure that represents the determinant of its Hessian. It is worth noticing that Theorems A-B-C hold purely under assumptions (1.1)-(1.5), hence they may be of independent interest. Next, we exploit Theorems A-B to analyze the inclusion (1.7) coupled with (1.1), and obtain partial $C^{1}$ regularity together with a precise description of the singular set in Theorem D (which extends [37] to general $\sigma$). Finally, we combine Theorems A-B-C-D with the very recent [34] to deduce partial $C^{1}$ regularity for solutions to the general equation (1.1), Theorem E.
In the course of the paper, it will be convenient to take into account three equivalent viewpoints: other than (1.1), it is better sometimes to look at solutions to differential inclusions, and sometimes at solutions to nonlinear Beltrami equations. While we defer the explanation of the latter to §2.2, where we will also thoroughly explain the interplay among the three, let us focus on the theory of differential inclusions in two dimensions in the next section. This will be useful to introduce our results in §1.2-1.3-1.4.
1.1. The differential inclusions point of view
Given a solution $u\in\operatorname{Lip}(B_{1})$ to (1.1), it is rather easy to construct a solution $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ to
@@ 1.1. The differential inclusions point of view
where $K\subset\mathbb{R}^{2\times 2}$ fulfills
@@
as soon as $G$ fulfills the monotonicity condition (1.5)-(1.6). The converse also holds: given a solution $w=(u,v)$ to (1.8) with $K$ enjoying property (1.9), then $u$ solves (1.1) for $G$ satisfying (1.5)-(1.6). The details of this translation is the content of Proposition 2.2 below.
In the theory of differential inclusions, ellipticity can be defined as in [54, §5], where a compact set $K$ is called elliptic if it fulfills (1.9), it is a manifold and its tangent space at every point has no rank-one connections. Recall that $A,B\in\mathbb{R}^{n\times m}$ are rank one-connected if $\operatorname{rank}(A-B)=1$. For consistency with the usual terminology in the world of PDEs, in this paper we refer to this requirement as uniform ellipticity. For example, the aforementioned $p$-Laplacian is an elliptic equation which does not admit elliptic linearizations at all points (hence is not uniformly elliptic) if $p\neq 2$. Hence, we give a weaker definition of ellipticity than [54], that a priori includes and extends the one carried by the PDE (1.1), where $G$ fulfills (1.5)-(1.6). While this is just a matter of terminology, it is crucial to underline the difference between our meaning of the word elliptic and the one of [54]: the fact that our elliptic sets may not admit elliptic tangent spaces at all points is the crux of the matter in Question (Q). In simple terms, our notion of ellipticity is equivalent to approximate rigidity of $K$:
Definition.
A compact set $K\subset\mathbb{R}^{n\times m}$ is called elliptic if, given any equi-Lipschitz sequence $(u_{k})_{k}$ on $B_{1}$,
@@
Equivalently, $K$ is elliptic if it only supports trivial homogeneous gradient Young measures. For a thorough introduction on Young measures, see for example [46].
The primary examples of elliptic sets are precisely those with property (1.9), compare [54, Theorem 1]:
@@
Showing that a set in $\mathbb{R}^{n\times m}$ is elliptic is, in general, very hard. In many situations, for instance in some convex integration schemes [47, 52], it is actually sufficient to show that the set is not elliptic by finding special subsets of matrices inside $K$: common choices of such special sets are rank-one connections and $T_{N}$ configurations, see [53, Definition 1]. If $K\subset\mathbb{R}^{2\times 2}$, we have a much clearer picture. In particular, from the deep works [53, 22] we have the following very simple way of deciding whether a given set in $\mathbb{R}^{2\times 2}$ is elliptic.
@@
We will add a few words on the essential range of the gradient of a map and its relation with Clarke’s generalized differential in §2.1.1.
@@
The statement of [32, Theorem 1] is more general than Theorem 1.3. It asserts that, for any $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$, the rank-one convex hull $Dw]^{\text{rc}}$ is connected. This implies Theorem 1.3 because $Dw]^{\text{rc}}=[Dw]$ if $Dw]\subset K$ for an elliptic set $K$. Indeed, the ellipticity assumption ensures that the quasiconvex hull of $[Dw]$, $Dw]^{\text{qc}}$, which fulfills $Dw]\subset[Dw]^{\text{rc}}\subset[Dw]^{\text{qc}}$, is equal to $[Dw]$, see [46, §4.4].

### v2

@@ 1. Introduction
The topic of this work is the regularity of two dimensional Lipschitz solutions to the elliptic equation
$\operatorname{div}(G(Du))=0\quad\text{ in }B_{1}\subset\mathbb{R}^{2}\,$, (1.1)
in the sense of distributions, for a general continuous field $G\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ which is strictly monotone:
$(G(b)-G(a),b-a)>0\quad\forall a\neq b\in\mathbb{R}^{2}\,.$ (1.2)
Specifically, we focus on the fundamental question:
[…]
Here, partially $C^{1}$ means $C^{1}$ in the complement of a closed Lebesgue-negligible set. For a continuous field $G$, (1.2) is, up to a change of sign, the weakest assumption under which Question (Q) is reasonable. Indeed, if $G(b)-G(a)\perp b-a$ for some $b\neq a$, then there exist Lipschitz solutions $u$ of (1.1) with wildly discontinuous gradient, namely $u(x)=\left(x,\frac{b+a}{2}\right)+f\left(\left(x,\frac{b-a}{2}\right)\right)$ for any $f\in\operatorname{Lip}(\mathbb{R})$ with $f^{\prime}\in\{\pm 1\}$ a.e..
Of particular and classical interest are gradient fields $G=Df$ with $f$ strictly convex: in that case,
$\operatorname{div}(Df(Du))=0\quad\text{ in }B_{1},$ (1.3)
is the Euler-Lagrange equation satisfied by minima of the functional
@@
It is well-known that solutions are smooth provided $f$ is smooth and uniformly elliptic, that is,
@@
In higher dimensions, this is due to the celebrated theorems of De Giorgi-Nash-Moser [18, 50, 47], while in two dimensions it has been known since the work of Morrey [46]. Many important and natural problems do not, however, enjoy the strong ellipticity property (1.5). A prominent example is that of the $p$-Laplace equation corresponding to $f(\xi)=|\xi|^{p}$. More generally, problems where (1.5) fails at a polynomial rate for $|\xi|\to 0$ and $|\xi|\to\infty$ have been studied extensively, see the review [44]. We are aware of very few results going beyond this type of degeneracy in higher dimensions: see [45, 43, 42], and [10, 9] for related regularity questions when $f$ is convex but not strictly convex.
Note that, due to the Lipschitz assumption, (Q) focuses on controlling oscillations of the gradient $Du$. A similar question could be asked about concentration effects, namely starting from a $W^{1,p}(B_{1})$ function. Some results in this direction, valid in all dimensions, can be found in [16, 17] and references therein. Recent counterexamples [11, 32] emphasize the necessity of careful assumptions on $u$ and $G$, but the Lipschitz property of $u$ can be inferred naturally in many situations, see e.g. [26, §12.4] and [8].
1.1. Previous two-dimensional results
Morrey’s original result [46] admits a first far-reaching generalization to $\delta$-monotone fields, which satisfy a more quantitative version of (1.2), see [2, §16]. This covers for instance the $p$-Laplace equation, but also many more degenerate cases: for instance (1.5) may fail on a large set of $\xi$’s as long as the eigenvalues of $D^{2}f(\xi)$ are comparable.
More generally and, to the best of our knowledge, for the first time, regularity results in the absence of any quantitative assumptions on the strict monotonicity (1.2) have been established by D. De Silva and O. Savin in [22]. For any strictly convex $f$, their results depend only on the closed sets $\mathcal{D}_{-}$ and $\mathcal{D}_{+}$ of values $\xi\in\mathbb{R}^{2}$ where ellipticity fails from below or from above:
[…]
The inequalities are in the sense of distributions with values into symmetric matrices. These definitions have natural extensions to any strictly monotone field $G$, see (9.1). It may be useful to have a picture of these sets in simple examples:
[…]
With this notation, the most general answer to (Q) known so far is, essentially, that
[…]
Quite remarkably, this only involves the degeneracy set $\mathcal{D}$ where ellipticity fails both from above and below. This regularity result is proved in [37] by building on the arguments introduced in [22], where $\mathcal{D}$ was assumed empty, or $\mathcal{D}_{-}$ finite. One can actually replace ‘finite’ in (1.6) by ‘at most countable’, with the same proof. Moreover, an example given in [37, Theorem 1.5] demonstrates that, under the mere assumption (1.2), Lipschitz solutions of (1.1) may have point singularities, so one cannot in general expect better than partial regularity in (Q). For different but related problems, some partial regularity results have been obtained in cases where the degeneracy set $\mathcal{D}$ may be one-dimensional: the boundary of a convex polygon for solutions of an obstacle problem [22] or a class of smooth curves in which $Du$ is constrained to lie [39].
1.2. Main results
Our first main result is valid for any strictly monotone continuous field $G$. We denote by $\mathcal{H}^{\alpha}$ the $\alpha$-dimensional Hausdorff measure.
Theorem 1.1.
Let $G\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ be continuous and strictly monotone (1.2), and $u\colon B_{1}\to\mathbb{R}$ a Lipschitz solution to (1.1). Then the nondifferentiability set
$\mathcal{S}_{u}=\{x\in B_{1}\colon u\text{ is not differentiable at }x\}$
satisfies $\mathcal{H}^{1}(\mathcal{S}_{u})=0$, and $Du$ is continuous on $B_{1}\setminus\mathcal{S}_{u}$.
Compared to a generic Lipschitz function, the gain of information provided by Theorem 1.1 is that:
  • the set of non-differentiability points is $\mathcal{H}^{1}$-negligible, instead of merely $\mathcal{L}^{2}$-negligible;
  • oscillations of $Du$ are controlled pointwise at all differentiability points, instead of merely in average at Lebesgue points of $Du$.
Thanks to the latter, partial $C^{1}$ regularity follows from Theorem 1.1 if one establishes that $\mathcal{S}_{u}$ is closed, or at least that its closure is $\mathcal{L}^{2}$-negligible. Our second main result achieves this, and more, under structural assumptions on the degeneracy set $\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}$ where ellipticity fails both from below and from above.
Theorem 1.2.
Let $G\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ be continuous and strictly monotone (1.2). Then, for any Lipschitz solution $u$ to (1.1), the singular set $\mathcal{S}_{u}\subset B_{1}$ is
  • locally finite, if the graph of $G$ over $\mathcal{D}$ is included in a finite disjoint union of $C^{1}$ curves in $\mathbb{R}^{2}\times\mathbb{R}^{2}$;
  • empty, if $\mathcal{D}$ is included in a countable disjoint union of boundaries of strictly convex sets.
We will actually prove a more general statement, see Theorem 9.1, but Theorem 1.2 conveys the main novelty of our methods, which is to deal with a large class of one-dimensional degeneracy sets $\mathcal{D}$, in contrast with the zero-dimensional assumption in (1.6). The first part of this result should be compared with the example found in [37, Theorem 1.5], of $u$ and $G$ solving (1.1)-(1.2) and $u$ having a point singularity. In that example, the graph of $G$ over $\mathcal{D}$ is a smooth curve. On one hand, it shows therefore the sharpness of the first part of Theorem 1.2. On the other hand, Theorem 1.2 shows the optimality of [37, Theorem 1.5] in the sense that point singularities are the only possible irregular phenomenon. The second part of Theorem 1.2 applies to a wide class of radial fields $G=Df$ with $f$ strictly convex and radial, a symmetry present in most physically relevant problems. More precisely, consider $f(x)=g(|x|)$ for some increasing and strictly convex $C^{1}$ function $g\colon[0,\infty)\to[0,\infty)$ such that $g(0)=g^{\prime}(0)=0$. Then we have $\xi\in\mathcal{D}$ if and only if $|\xi|\in D_{g}$, where
[…]
Informally, one can think of $D_{g}$ as the set of points $t\geq 0$ near which $g^{\prime\prime}$ can be both arbitrarily large and arbitrarily small. If $D_{g}$ is totally disconnected, then all connected components of $\mathcal{D}$ are either circles or the origin, and we deduce that any Lipschitz minimizer of (1.4) is $C^{1}$. This is entirely new: in the context of radial functionals, the available techniques so far could only show such regularity under the assumption that $\mathcal{D}$ is precisely the origin. We conclude this introduction by explaining the strategy and the tools used to show the above results, other than giving an overview of the structure of the paper.
1.3. Strategy of Theorem 1.1
Given some uniform monotonicity in (1.2), a classical first step is to obtain information on $Du$ by choosing suitable test functions and obtaining integral estimates. This step is present for instance in the aforementioned works [46, 22, 2, 39], and, more generally, all regularity results we are aware of require quantitative forms of the monotonicity (1.2) at least in suitable open sets. However, to solve Question (Q) in full generality, one needs to deal with arbitrarily degenerate monotonicity on any open set. For instance, the degeneracy set $\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}$ can fill the whole plane, see Remark 9.1.
Thus, we need to rely on different and novel tools: the main point in the proof of Theorem 1.1 is an $L^{\infty}$ to $W^{1,\infty}$ estimate inspired by the beautiful separation argument of [24]. It states that pointwise gradient localization into a given convex region is stable with respect to the $L^{\infty}$ topology, see Theorem 4.1. Since linear functions solve (1.1) and their gradient is localized at a single value, this implies a flatness improvement property: if a solution $u$ of (1.1) with $|Du|\leq L$ is uniformly close to a linear function $\ell_{\xi}(x)=(\xi,x)$, then $Du$ is uniformly close to $\xi$ in a smaller ball. More precisely, $\forall\xi\in\overline{B_{L}},$
$\forall\varepsilon>0,\,\exists\delta>0\colon\quad\|u-\ell_{\xi}\|_{L^{\infty}(B_{1})}\leq\delta\quad\Rightarrow\quad\|Du-\xi\|_{L^{\infty}(B_{\delta})}\leq\epsilon\,.$ (1.7)
Analogous flatness results are at the heart of [10, 9, 45], but with the fundamental difference that they require $\xi$ to be an elliptic value, that is, $\xi\notin\mathcal{D}_{-}\cup\mathcal{D}_{+}$. Similar arguments can be found also in [22, 37], provided that $\xi\notin\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}$, but observe that this might be an empty statement in our general context, since $\mathcal{D}$ could be the whole plane. Here, instead, we can apply (1.7) for any $\xi\in\mathbb{R}^{2}$, and infer in particular that if $Du$ has vanishing mean oscillation (VMO) at $x\in B_{1}$, then $u$ is differentiable at $x$, and $Du$ has vanishing pointwise oscillation at $x$. This implies Theorem 1.1, provided we show that non VMO points of $Du$ are $\mathcal{H}^{1}$-negligible. This last fact follows from the strict monotonicity of $G$, and is actually valid in any dimension (with $\mathcal{H}^{1}$ replaced by $\mathcal{H}^{n-1}$), see Theorem 5.1.
Note that, from Theorem 1.1, a classical way to show that $\mathcal{S}_{u}$ is closed, hence prove partial regularity, would be to combine it with an $\varepsilon$-regularity result:
$\inf_{\xi\in\overline{B_{L}}}\|Du-\xi\|_{L^{\infty}(B_{1})}\leq\varepsilon\quad\Longrightarrow\quad Du\text{ continuous in }B_{\varepsilon}\,$, (1.8)
for an $\varepsilon>0$ depending only on $G$ and $L$. However, there exist continuous strictly monotone fields for which such an $\varepsilon>0$ cannot exist, see Remark 9.1, so the full answer to (Q) would have to follow a different route.
1.4. Strategy of Theorem 1.2
The starting point is the recent and remarkable localization result of [36], which states that any blowup limit
$u_{\infty}(x)=\lim_{r_{j}\to 0}\frac{u(x_{0}+r_{j}x)-u(x_{0})}{r_{j}}\,$,
is either linear or satisfies $Du_{\infty}\in\mathcal{D}$ a.e. in $\mathbb{R}^{2}$. At a singular point, we are in the latter situation, and blowup limits are solutions of the constrained problem
$\operatorname{div}G(Du_{\infty})=0,\text{ and }Du_{\infty}\in\mathcal{D}\,.$ (1.9)
We use this constrained problem to show that blowup limits are one-homogeneous with a single singularity at the origin: standard arguments then imply that $\mathcal{S}_{u}$ is locally finite. It is worth emphasizing that homogeneity of blowup limits is typically obtained in elliptic problems through monotonicity formulas, which are unavailable in our problem. Here, in contrast, we obtain this homogeneity relying on:
  1. (1)
    hyperbolic regularization effects hidden in (1.9) when $\mathcal{D}$ is a curve, see Theorem 8.1;
  2. (2)
    the observation of independent interest that general solutions to (1.1) have a well-defined Monge-Ampère measure $\det(D^{2}u)\leq 0$, which concentrates at the origin for blowup limits, see Theorem 7.1 and Corollary 7.
We will comment further on the second point in §7, but let us be more specific about the first point, namely the hyperbolic regularization effect, which yields a strong generalization of the main result of [39] about the regularity of differential inclusions into curves. The starting observation is that, when $\mathcal{D}$ is a curve, the degenerate elliptic problem (1.9) becomes a hyperbolic system, and $C^{1}$ solutions have constant gradients along characteristic lines. A typical feature of hyperbolic problems is that weak solutions can be extremely wild, but regularization effects are available for a subclass of weak solutions. Here we manage to exploit such effects and eventually obtain an $\varepsilon$-regularity result (1.8) for solutions of (1.9). An observation from [39] relating (1.9) to zero-energy states of a phase transition model [27] then enables us to deduce that the singular set $\mathcal{S}_{u_{\infty}}$ is locally finite. For entire solutions of (1.9), such as blowup limits, the constancy along characteristics outside $\mathcal{S}_{u_{\infty}}$ further enforces a quite rigid structure, which we can finally combine with the concentration property of the Monge-Ampère measure to conclude that blowup limits are one-homogeneous.
1.5. The differential inclusions approach and nonlinear Beltrami equations
Ubiquitous in our approach is the correspondence between Lipschitz solutions of (1.1) and solutions to differential inclusions. In order to make a short but comprehensive overview of the subject, we postpone a thorough explanation of this theory to §2. Alongside differential inclusions, we will also exploit the equivalent viewpoint of nonlinear Beltrami systems, which will be introduced in §3.2. There, we will explain in detail the (well-known) interplay and equivalence among these three viewpoints. It is important to notice that, on one hand, the theory of $2\times 2$ elliptic differential inclusions is extremely rich, and this wealth of results and techniques will be essential in our paper. Conversely, our results, especially Theorem 1.1, represent an important step towards a complete understanding of elliptic differential inclusions in $\mathbb{R}^{2\times 2}$, see Question (Q:DI) below.
Structure of the paper
We begin by introducing differential inclusions in §2. In §3, we explain the notation and the well-known, but crucial, results mentioned in §1.5. §45 contain our first two main Theorems 4.1-5.1 that, combined, yield Theorem 1.1, proved in §6. In §7 we introduce the Monge-Ampère measure associated to a solution of the elliptic PDE (1.1). §8 contains a sharp result on inclusions into elliptic curves, Theorem 9.1. All of these results will be used in the proof of Theorem 1.2, contained in §9. Some sections are lengthier than others: in that case we will describe the section at the start of it.
Acknowledgments
XL is supported by the ANR project ANR-22-CE40-0006. We thank L. De Rosa and A. Guerra for highly valuable comments on the presentation.
2. The differential inclusions point of view
Before explaining some elements from the theory of differential inclusions and how they relate to our paper, let us set a notation which will be useful throughout: we express condition (1.2) equivalently as
$(G(b)-G(a),b-a)\geq\sigma(|a-b|)\quad\forall a,b\in\mathbb{R}^{2}\,$, (2.1)
@@ 2. The differential inclusions point of view
$\sigma:[0,+\infty)\to[0,+\infty)\text{ is a continuous, strictly increasing, convex function.}$ (2.2)
Now, given a solution $u\in\operatorname{Lip}(B_{1})$ to (1.1), one can construct a solution $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ to
@@
where $K\subset\mathbb{R}^{2\times 2}$ fulfills, for a possibly different $\sigma$ still fulfilling (2.2),
@@
as soon as $G$ fulfills the monotonicity conditions (2.1)-(2.2). The converse also holds: given a solution $w=(u,v)$ to (2.3) with $K$ enjoying properties (2.4)-(2.2), then $u$ solves (1.1) for $G$ satisfying (2.1)-(2.2), see Proposition 3.2 below. In the following, for brevity we will often say that $G$ (or $K$) fulfills (2.1) (or (2.4)) to say that $G$ (or $K$) fulfills (2.1) (or (2.4)) for $\sigma$ fulfilling (2.2).
For differential inclusions, ellipticity can be defined as in [56, §5], where a compact set $K$ is called elliptic if it fulfills (2.4), it is a manifold and its tangent space at every point has no rank-one connections. Recall that $A,B\in\mathbb{R}^{n\times m}$ are rank one-connected if $\operatorname{rank}(A-B)=1$. For consistency with the usual terminology in the world of PDEs, in this paper we refer to this requirement as uniform ellipticity. For example, the $p$-Laplacian is an elliptic equation which does not admit elliptic linearizations at all points (hence is not uniformly elliptic) if $p\neq 2$. Thus, we give a weaker definition of ellipticity than [56], that a priori includes and extends the one carried by the PDE (1.1), where $G$ fulfills (2.1)-(2.2). While this is just a matter of terminology, it is crucial to underline the difference between our meaning of the word elliptic and the one of [56]: the fact that our elliptic sets may not admit elliptic tangent spaces at all points is the crux of the matter in Question (Q). In simple terms, our notion of ellipticity is equivalent to approximate rigidity of $K$:
Definition 2.1.
A set $K\subset\mathbb{R}^{n\times m}$ is called elliptic if, given any equi-Lipschitz sequence $(u_{k})_{k}$ on $B_{1}$,
@@
Equivalently, a compact set $K$ is elliptic if it only supports trivial homogeneous gradient Young measures. For a thorough introduction on Young measures, see for example [48].
The primary examples of elliptic sets are precisely those with property (2.4), compare [56, Theorem 1]:
@@
Showing that a set in $\mathbb{R}^{n\times m}$ is elliptic is, in general, very hard. In many situations, for instance in some convex integration schemes [49, 54], it is actually sufficient to show that the set is not elliptic by finding special subsets of matrices inside $K$: common choices of such special sets are rank-one connections and $T_{N}$ configurations, see [55, Definition 1]. If $K\subset\mathbb{R}^{2\times 2}$, we have a much clearer picture. In particular, from the deep works [55, 24] we have the following very simple way of deciding whether a given set in $\mathbb{R}^{2\times 2}$ is elliptic.
@@
We will add a few words on the essential range of the gradient of a map in §3.1.1.
@@
The statement of [34, Theorem 1] is more general than Theorem 2.4. It asserts that, for any $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$, the rank-one convex hull $Dw]^{\text{rc}}$ is connected. This implies Theorem 2.4 because $Dw]^{\text{rc}}=[Dw]$ if $Dw]\subset K$ for an elliptic set $K$. Indeed, the ellipticity assumption ensures that the quasiconvex hull of $[Dw]$, $Dw]^{\text{qc}}$, which fulfills $Dw]\subset[Dw]^{\text{rc}}\subset[Dw]^{\text{qc}}$, is equal to $[Dw]$, see [48, §4.4].